OpenStudy (anonymous):

Solve the equation. 9x^2 - 4 = 0

OpenStudy (angela210793):

9x^2=4 x^2=4/9 x=+-sqr4/9=+-2/3

OpenStudy (anonymous):

9x^2=4 x=+-2/3

OpenStudy (anonymous):

9x^2=4 9x=sqrt4 9x=4 x=4/9 x=2/3

OpenStudy (angela210793):

if u sqr both parts of the eq u'd have 3x=+-2 x=+-2/3

OpenStudy (anonymous):

=(3x+2)*(3x-2) x=+or-2/3

OpenStudy (anonymous):

\[9x^2 - 4 = 0 \rightarrow9x^2=4 \rightarrow x^2=\frac{4}{9} \rightarrow x=\sqrt{\frac{4}{9}}\]